OpenStudy (anonymous):

Use Gaussian elimination to find the complete solution to the system of equations, or state that none exists. x + y + z = 9 2x - 3y + 4z = 7 x - 4y + 3z = -2

OpenStudy (amistre64):

form the coefficient matrix and row reduce to echelon form

OpenStudy (amistre64):

the last column will produce the solution

OpenStudy (anonymous):

ok let me try thanks

OpenStudy (amistre64):

if i didnt hit a wrong key .... looks like a row of zeros on the bottom for "infinite" solutions

OpenStudy (amistre64):

lets test the determinant -(-3-16+6) 1 1 1 1 1 2 -3 4 2 -3 1 -4 3 1 -4 +(-9+4-8) (-9+4-8) - (-3-16+6) (-17+4) - (-3-10) (-13) - (-13) -13 + 13 = 0

OpenStudy (anonymous):

here is what i came up with 7z/5+ 34/5+2z/5+ 11/5z) looks like were off

OpenStudy (anonymous):

I think it might be negitive im confused

OpenStudy (amistre64):

1x + 1y + 1z = 9 2x - 3y + 4z = 7 1x - 4y + 3z = -2 \[\begin{pmatrix}1&1&1\\2&-3&4\\1&-4&3\\\end{pmatrix}~\begin{pmatrix}x\\y\\z\end{pmatrix}=\begin{pmatrix}9\\7\\-2\end{pmatrix}\] \[\begin{pmatrix}1&1&1&|&9\\2&-3&4&|&7\\1&-4&3&|&-2\end{pmatrix}\to\begin{pmatrix}1&0&1.4&|&6.8\\0&1&-.4&|&2.2\\0&0&0&|&0\end{pmatrix}\] \[x=6.8-1.4z\] \[x=2.2+0.4z\] \[z=~0~+1.0z\]

OpenStudy (amistre64):

that was my ti83, the wolf does this rref{{1,1,1,9},{2,-3,4,7},{1,-4,3,-2}} http://www.wolframalpha.com/input/?i=rref%7B%7B1%2C1%2C1%2C9%7D%2C%7B2%2C-3%2C4%2C7%7D%2C%7B1%2C-4%2C3%2C-2%7D%7D

OpenStudy (amistre64):

either way, we get an under row of all zeros, which tells me that 0 is always equal to 0, so infinite solutions

OpenStudy (anonymous):

thank you very much!!

OpenStudy (amistre64):

good luck :)